# Branches, Assemble!  Multi-Branch Cooperation Network for Large-Scale Click-Through Rate Prediction at Taobao

### v1

@@ 4.2.2. Online Production Deployment
We further evaluate the method through online A/B testing in image2product search of Alibaba Taobao system. In this experiment, the training data is our Pailitao-24month dataset. The baseline is our previously online serving model. Online evaluation metrics are real CTR, deal number by query, deal number by user and GMV. Results in Table 4 illustrate that our MBCnet has an absolute 0.09 point CTR increase, a relative 1.49% deal growth and a relative 1.62% GMV increase. In Figure 3, we also present the CTR improvements compared to online base model across query categories and user groups. The figure clearly shows that MBCnet achieves significant improvements from multiple evaluation perspectives, highlighting its effectiveness in our production system. Since August 2024, MBCnet has been fully deployed online, serving hundreds of millions of customers.
@@ 4.3. Why Using Multi-Branch Cooperation
In this part, we conduct an experiment to demonstrate why using multi-branch cooperation in large-scale CTR prediction. In particular, we show some learning dynamics of different branches during training. Moreover, we compare the feature distribution before logit layer across branches with t-SNE (van der Maaten and Hinton, 2008). The results are summarized in Figure 4, where “EFGC”, “Deep”, “CrossNet” are three branches and “MBC” is the fused one.
From Figure 4 (a)-(c), we see that: 1) The fused ”MBC” model consistently achieves lower training log loss and validation log loss, along with a higher validation AUC throughout the training process. This indicates that a single branch has limited learning capacity, whereas MBCnet enhances its learning ability by cooperating multiple branches. As shown in Figure 4 (d), the output features from ”EFGC,” ”Deep,” and ”CrossNet” branches are distinct from one another. In contrast, the fused ”MBC” features are more evenly distributed across the feature space. This demonstrates that each branch has its own modeling patterns, and MBCnet can integrate their diverse capabilities. These findings emphasize the value of utilizing collaboration among branches to model the complex distributions of user behaviors. More experiments about why using multi-branch cooperation are given in Appendix A.1 and A.2.
@@ 4.4. How to do Multi-Branch Cooperation
In Table 5, we show that violating our proposed scheme leads to a decline in model performance. Principle 1 is particularly crucial because it dictates the selection of knowledge to be transferred among branches. Principle 2 also affects performance by ensuring branch diversity and encouraging the model to discover different patterns. Additionally, Figure 5 reveals that both the co-teaching loss and the moderate differentiation loss converge, while the validation AUC steadily increases. In the converged state, the difference between equivalent transformation matrix $W$ and $I_{d}$ remains non-zero, demonstrating that the branches maintain distinct discrepancies to explore diverse patterns effectively.
@@ 4.5. Study of Different Model Parts
To evaluate the impact of each model component, we performed an ablation study by systematically removing specific modules, as detailed in Table 6. The results indicate a clear decline in performance on both datasets when any component is omitted. For instance, excluding the EFGC branch leads to a 0.94% decrease in AUC on Pailitao-24month. This result highlights the importance of our EFGC branch. The mechanism in EFGC highlights feature crossing within feature groups, which encourages the model to capture specific patterns and meanwhile reduces redundant crossing features. Additionally, removing the cooperation scheme, i.e., “w/o $\mathcal{L}_{\text{BCT}}$, w/o $\mathcal{L}_{\text{MDR}}$”, results in a 0.79% reduction in AUC on Pailitao-12month. These outcomes demonstrate that each component positively contributes to the overall performance of our model.
@@ 5. Conclusion and Future Work
In this paper, we introduced Multi-Branch Cooperation Network (MBCnet) for enhancing CTR prediction in large-scale online services. MBCnet effectively integrates three distinct branches (i.e., EFGC, low rank CrossNet and Deep Net) in capturing complex feature interactions. The proposed cooperation scheme, based on the principles of branch co-teaching and moderate differentiation, facilitates meaningful collaboration between branches. It also enables the model to leverage diverse patterns and improve overall prediction accuracy. Our extensive experiments, including an online A/B test, demonstrated significant improvements in CTR, deal growth, and GMV, validating its effectiveness and scalability in real-world applications.
@@
In future work, we plan to incorporate more powerful branches to further improve the model’s performance. Meanwhile, we use loss values to classify branches as strong or weak on particular samples. This method may produce unreliable loss measurements during the early training stage, potentially limiting the model’s learning ability. We will develop more effective techniques for identifying strong and weak branches to achieve better prediction performance.
Appendix A Additional Experiments
A.1. Branch Difference at Neuron Level
@@
We design an experiment to illustrate the branch difference at neuron level. In particular, MBCnet integrates latent features from multiple branches to execute the final CTR prediction task. These latent features are derived from the outputs of neurons, as depicted in Figure 6 (a). Each dimension of a feature value reflects the strength of a specific high-order feature. To analyze these feature strengths for each branch, we normalize the neuron outputs using softmax function [^5] Our network has an output dimension of 128. For clarity, we randomly select 30 neuron outputs to illustrate the results.. The normalized distributions are shown in Figure 6 (b) [^6] Notice that the output features from different branches are in the same space within our network design, enabling direct comparison of neuron values across branches..
@@
From the figure, we have the following observations: 1) Different branches have close outputs at most neurons while quite different values at few neurons. This indicates different branches have many common learned patterns and several specific patterns, which aligns with our moderate differentiation principle. The common patterns ensure branches are learned towards the same learning objective, while the specific patterns encourage the model to develop complementary functionalities, benefiting the final prediction task. 2) We also observe that “EFGC” branch exhibits greater consistency with “CrossNet” branch compared to “Deep” branch. One possible reason for this may be that the “EFGC” branch engages in feature crossing among expert-selected features, which can be regarded as one kind of explicit feature crossing in “CrossNet”.
@@
A.2. Branch Difference at Category Level
@@
A.3. Adding More Branches into MBCnet
@@
A.4. Hyper-Parameter Sensitivity

### v2

@@ 4.2.2. Online Production Deployment
We further evaluate the method through online A/B testing in image2product search of Alibaba Taobao system. In this experiment, the training data is our Pailitao-24month dataset. The baseline is our previously online serving model (DCNv2). Online evaluation metrics are real CTR, deal number by query, deal number by user and GMV. Table 4 shows that our MBCnet has an absolute 0.09 point CTR increase, a relative 1.49% deal growth and a relative 1.62% GMV increase. In Figure 3, we also present the CTR improvements compared to online base model across query categories and user groups. The figure clearly shows that MBCnet achieves significant improvements from multiple evaluation perspectives, highlighting its effectiveness in our production system. Since August 2024, MBCnet has been fully deployed online, serving hundreds of millions of customers.
Meanwhile, different components of MBCnet work in a parallel style so that we can use parallel computation to accelerate training and inference. When conducting the online A/B test, the online base model is DCNv2 (i.e. CrossNet and Deep branch), and MBCnet has additional EFGC branch for inference. The EFGC branch is simple MLP and has a 2.2 million increased parameters, which is small for industrial CTR models. In our experiments, the real-time latency only has 1ms increase and the GPU utility only has 2% increase, which are negligible in practice. Detailed time complexity analysis is given in Appendix C.
@@ 4.3. Why Using Multi-Branch Cooperation
In this part, we conduct an experiment to demonstrate why using multi-branch cooperation in our industrial business. In particular, we show some learning dynamics of different branches during training. Moreover, we compare the feature distribution before logit layer across branches with t-SNE (van der Maaten and Hinton, 2008). The results are summarized in Figure 4, where “EFGC”, “Deep”, “CrossNet” are three branches and “MBC” is the fused one.
From Figure 4 (a)-(c), we see that: 1) The fused ”MBC” model consistently achieves lower training log loss and validation log loss, along with a higher validation AUC throughout the training process. This indicates that a single branch has limited learning capacity, whereas MBCnet enhances its learning ability by cooperating multiple branches. 2) As shown in Figure 4 (d), the output features from ”EFGC,” ”Deep,” and ”CrossNet” branches are distinct from one another. In contrast, the fused ”MBC” features are more evenly distributed across the feature space. In fact, we have also observed the output features across branches are more entangled and hard to be identified in early training stage, while the features tend to cluster to different subspaces and the fused “MBC” is becoming more evenly distributed across the feature space in later training epochs. These cooperative dynamics demonstrate that each branch has its own modeling patterns, and MBCnet can gradually integrate their diverse capabilities. These emphasize the value of utilizing collaboration among branches to model the complex distributions of user behaviors. More experiments about why using multi-branch cooperation are given in Appendix B.1 and B.2.
@@ 4.4. How to do Multi-Branch Cooperation
In Table 5, we show that violating our proposed scheme leads to a decline in model performance. Principle 1 is particularly crucial because it dictates the selection of knowledge to be transferred among branches. Principle 2 also affects performance by ensuring branch diversity and encouraging the model to discover different patterns. Additionally, Figure 5 reveals that both the co-teaching loss and the moderate differentiation loss converge, while the validation AUC steadily increases. In the converged state shown in Figure 5 (c), the difference between equivalent transformation matrix $W$ and $I_{d}$ remains non-zero, demonstrating that the branches maintain distinct discrepancies to explore diverse patterns.
@@ 4.5. Study of Different Model Parts
To evaluate the impact of each model component, we performed an ablation study by systematically removing specific modules, as detailed in Table 6. The results indicate a clear decline in performance on both datasets when any component is removed. For instance, excluding the EFGC branch leads to a 0.94% decrease in AUC on Pailitao-24month. This result highlights the importance of our EFGC branch. The mechanism in EFGC highlights feature crossing within feature groups, which encourages the model to capture specific patterns and meanwhile reduces redundant crossing features. Additionally, removing the cooperation scheme, i.e., “w/o $\mathcal{L}_{\text{BCT}}$, w/o $\mathcal{L}_{\text{MDR}}$”, results in a 0.79% reduction in AUC on Pailitao-12month. These outcomes demonstrate that each component positively contributes to the overall performance of our model. More experiments and analysis are provided in the Appendix.
@@ 5. Conclusion and Future Work
In this paper, we introduced Multi-Branch Cooperation Network (MBCnet) for enhanced CTR prediction. MBCnet effectively integrates three distinct branches (i.e., EFGC, low rank CrossNet and Deep Net) in capturing complex feature interactions. The proposed cooperation scheme, based on the principles of branch co-teaching and moderate differentiation, facilitates meaningful collaboration between branches. It also enables the model to explore diverse feature interaction patterns and improve the overall prediction accuracy. Our extensive experiments, including an online A/B test, demonstrated significant improvements in CTR, deal growth, and GMV, validating its effectiveness and scalability in real-world applications.
@@
Currently we simply use the common average pooling and use later MLP layer to implicitly combine the outputs of branches. We believe that a better way to fuse the branch outputs could lead to better performance. Meanwhile, we use loss values to classify branches as strong or weak on particular samples. This may produce unreliable loss measurements during the early training stage, potentially limiting the model’s learning ability. We will study these issues in future work.
Appendix A About Low-Rank Cross Net
The low-rank cross network is from DCNv2 (Wang et al., 2021), and is a low-rank and computation-efficient version of the original cross network in DCN (Wang et al., 2017a). The original cross network explicitly learn high-order feature interactions through defined equation $x_{l+1}=x_{0}x_{l}^{T}w_{l}+b_{l}+x_{l}$, where $x_{l},x_{l+1}\in\mathbb{R}^{d}$ are column vectors denoting the outputs from $l$-th and $(l+1)$-th cross layers, respectively. $w_{l},b_{l}\in\mathbb{R}^{d}$ are the weight and bias parameters of the $l$-th layer. The special structure of the cross network causes the degree of cross features to grow with layer depth. The highest polynomial degree (in terms of input $x_{0}$ for an $l$ layer cross network is $l+1$. In fact, the cross network comprises all the cross terms. Refer to (Wang et al., 2021) for detailed proof.
This original cross network involves a dimension-consistent feature crossing in feature space, which would cause large computation when the dimension of inputs is high. Therefore, the low-rank cross network (Wang et al., 2021) aims to accelerate the feature interaction in original cross network. Specifically, it maps the feature into low-rank space and then maps it back to the original feature dimension with mixture of experts.
Figure 6. (a) is an example to illustrate the structure of neuron outputs before logit layer. (b) shows the output distribution of those neurons on Pailitao-12month.
Appendix B Additional Experiments
B.1. Branch Difference at Neuron Level
@@
We design an experiment to illustrate the branch difference at neuron level. In particular, MBCnet integrates latent features from multiple branches to execute the final CTR prediction task. These latent features are derived from the outputs of neurons, as depicted in Figure 6 (a). Each dimension of a feature value reflects the strength of a specific high-order feature. To analyze these feature strengths for each branch, we normalize the neuron outputs using softmax function [^5] Our network has an output dimension of 128. For clarity, we randomly select 30 neuron outputs to illustrate the results.. The normalized distributions are shown in Figure 6 (b) [^6] Notice that the output features from different branches are in the same space in our network design, enabling direct comparison of neuron values across branches..
@@
From the figure, we have the following observations: 1) Different branches have close outputs at most neurons while quite different values at few neurons. This indicates different branches have many common learned patterns and several specific patterns, which aligns with our moderate differentiation principle. The common patterns ensure branches are learned towards the same learning objective, while the specific patterns encourage the model to develop complementary functionalities, benefiting the final prediction task. 2) We also observe that “EFGC” branch exhibits greater consistency with “CrossNet” branch compared to “Deep” branch. One possible reason for this may be that the “EFGC” branch engages in feature crossing by domain-driven knowledge, which can be regarded as one kind of explicit feature crossing in “CrossNet”.
@@
B.2. Branch Difference at Category Level
@@
B.3. Adding More Branches into MBCnet
@@
B.4. Hyper-Parameter Sensitivity
@@
Table 7. The AUC comparison of different models under different data density on Pailitao-12month. The listed density ratio is relative to the original train set.
Density ratio 20% 40% 60% 80% 100%
DNN 0.5132 0.5885 0.6770 0.7223 0.7423
DCNv2 0.5200 0.5929 0.6797 0.7262 0.7461
MBCnet 0.5277 0.6015 0.6890 0.7319 0.7522
B.5. Performance Under Different Data Sparsity
We also conducted performance comparison under different data density. Specifically, we keep the validation and test set fixed, and mask a certain ratio of positive samples. The AUC results of some methods on Pailitao-12month are given in Table 7. The result shows that MBCnet consistently achieves better performance even in more sparse cases.
Appendix C Time Complexity Analysis
In Table 8, we have made a complexity analysis of different branches/components in MBCnet. The time complexity of EFGC branch is $\mathcal{O}(N_{g}L_{\text{EFGC}}d^{2})$, where $N_{g}$ is the number of feature groups, $L_{\text{EFGC}}$ is the number of EFGC layers and $d$ denotes the latent dimension. The time complexity of CrossNet branch is $\mathcal{O}(L_{\text{CrossNet}}N_{\text{experts}}Fd)$, where $L_{\text{CrossNet}}$ is the number of CrossNet layers, $N_{\text{experts}}$ denotes the number of experts and $F$ is the input feature dimension. For Deep branch, the time complexity is $\mathcal{O}(L_{\text{Deep}}d^{2})$, where $L_{\text{Deep}}$ is the number of mlp layers. Considering the branch fusion in Eq.15 is average pooling and the complexity is negligible, so the complexity of MBCnet can be taken as the summation of three branches.
Fortunately, different components of MBCnet works in a parallel style so that we can use parallel computation to accelerate training and inference. In our experiments, the real-time latency only has 1ms increase and the GPU utility only has 2% increase, which are negligible in practice.
Table 8. Time complexity analysis of different model components.
Method EFGC CrossNet Deep
Complexity $\mathcal{O}(N_{g}L_{\text{EFGC}}d^{2})$ $\mathcal{O}(L_{\text{CrossNet}}N_{\text{experts}}Fd)$ $\mathcal{O}(L_{\text{Deep}}d^{2})$
Ilustration
$N_{g}$: number of feature groups
$L_{\text{EFGC}}$: number of EFGC layers
$d$: latent dimension.
$L_{\text{CrossNet}}$: number of CrossNet layers
$N_{\text{experts}}$: number of experts
$F$: input feature dimension.
$L_{\text{Deep}}$: number of mlp layers
Appendix D Core Codes of MBCnet
To better illustrate the scheme of our cooperation idea, we provide the Tensorflow codes in Listing 1. The loss_op function defines the entrance of our training objective.
Listing 1:  Loss Code of MBCnet in Tensorflow Style
import tensorflow as tf
# binary cross entropy loss
def bce_loss_op(self, logit, label, reduction=’mean’):
if reduction==’mean’:
bce_loss = tf.reduce_mean(tf.nn.sigmoid_cross_entropy_with_logits(logits=logit,labels=label))
elif reduction==’sum’:
bce_loss = tf.reduce_sum(tf.nn.sigmoid_cross_entropy_with_logits(logits=logit,labels=label))
elif reduction==’none’:
bce_loss = tf.nn.sigmoid_cross_entropy_with_logits(logits=logit,labels=label)
else:
print(’Notsupportedreduction!’)
raise Exception
return bce_loss
# branch co-teaching loss
def BCT_loss_op(self, logits_list, labels):
losses_list = []
for logits in logits_list:
loss = self.bce_loss_op(logits, labels, reduction=’none’)
losses_list.append(loss)
BCT_loss = 0.0
n_sample = 1e-8
for i in range(len(logits_list)):
for j in range(i,len(logits_list)):
if i!=j:
# mask of disagreement data
###### simplified version
mask_i = tf.logical_and(tf.reshape(tf.cast(losses_list[i] < -tf.log(0.5), tf.bool), [-1]), tf.reshape(tf.cast(losses_list[j] > -tf.log(0.5), tf.bool),[-1]))
mask_j = tf.logical_and(tf.reshape(tf.cast(losses_list[j] < -tf.log(0.5), tf.bool), [-1]),tf.reshape(tf.cast(losses_list[i] > -tf.log(0.5), tf.bool),[-1]))
# label supervision from strong branch
BCT_loss += tf.cond(tf.reduce_sum(tf.cast(mask_j,tf.float32))>0, lambda:self.bce_loss_op(tf.boolean_mask(logits_list[i],mask_j),tf.stop_gradient(tf.boolean_mask(tf.sigmoid(logits_list[j]),mask_j)),reduction=’sum’), lambda:tf.zeros_like(self.reg_loss))
BCT_loss += tf.cond(tf.reduce_sum(tf.cast(mask_i,tf.float32))>0, lambda:self.bce_loss_op(tf.boolean_mask(logits_list[j],mask_i),tf.stop_gradient(tf.boolean_mask(tf.sigmoid(logits_list[i]),mask_i)),reduction=’sum’), lambda:tf.zeros_like(self.reg_loss))
n_sample += tf.reduce_sum(tf.cast(mask_j,tf.float32)) + tf.reduce_sum(tf.cast(mask_i,tf.float32))
BCT_loss = BCT_loss/n_sample
return BCT_loss
# moderate differentiation loss
def MDR_loss_op(self, feats_list):
mdr_loss_list = []
hidden_dim = feats_list[0].get_shape().as_list()[1]
for i in range(len(feats)):
for j in range(len(feats)):
if i!=j:
feat_i, feat_j = feats[i], feats[j]
# orthogonal transformation constraint
[…]
l2_distance = tf.norm(mapped_feat_i_to_j-feat_j, ord=’euclidean’, axis=1) + tf.norm(mapped_feat_i_to_j_to_i-feat_i, ord=’euclidean’, axis=1)
mdr_loss_list.append(l2_distance)
mdr_loss = tf.reduce_mean(tf.concat(mdr_loss_list, 0))
return mdr_loss
# The overall loss op
def loss_op(self, branch_logits_list, branch_top_feats_list, fused_logits, labels):
␣␣␣␣branch_logits_list:alistofbranchoutputlogits,[efgc_logits,crossnetv2_logits,dnn_logits],eachoneis[B,1]
␣␣␣␣branch_top_feats_list:alistofbranchoutputfeaturesbeforelogitlayer,[efgc_top_feat,crossnetv2_top_feat,dnn_top_feat],eachoneis[B,1]
␣␣␣␣fused_logits:thefusedbranchlogits,[B,1]
␣␣␣␣labels:thelabelsofsamples,[B,1]
efgc_logits, crossnetv2_logits, dnn_logits = branch_logits_list
with tf.name_scope("CTR_Loss_Op"):
# EFGC bce loss
efgc_bce_loss = self.bce_loss_op(efgc_logits)
# CrossNetV2 bce loss
crossnetv2_bce_loss = self.bce_loss_op(crossnetv2_logits)
# DNN bce loss
dnn_bce_loss = self.bce_loss_op(dnn_logits)
# fused bce loss
fused_bce_loss = self.bce_loss_op(fused_logits)
# overall bce loss
self.bce_loss = efgc_bce_loss + crossnetv2_bce_loss + dnn_bce_loss + fused_bce_loss
# multi-branch cooperation loss
self.BCT_loss = self.BCT_loss_op(branch_logits_list, labels)
self.MDR_loss = self.MDR_loss_op(branch_top_feats_list)
# final optimization loss
self.loss = self.bce_loss + self.alpha * self.BCT_loss + self.beta * self.MDR_loss